OpenStudy (liv33):

Solve for t: 5/13t = -9

TheSmartOne (thesmartone):

multiply both sides by the reciprical of 5/13 the recriprocal of a/b is b/a And I'm assuming the question looks like: \(\Large \frac{5}{13} t = -9\)

OpenStudy (liv33):

Ok (and, yes, that's how the questions looks)

TheSmartOne (thesmartone):

great, so what is the reciprocal?

OpenStudy (liv33):

13/5

TheSmartOne (thesmartone):

now, multiply it on both sides 13/5 * 5/13 * t = -9 * 13/5 multiplying by the reciprical helps us isolate the variable t, can you simplify that ^ ?

OpenStudy (liv33):

t = -23.4

TheSmartOne (thesmartone):

correct, did they want the final answer as a decimal or a fraction? :)

OpenStudy (liv33):

Either one, I assume. They only ask that I solve it, not to put the answer in one form or the other

TheSmartOne (thesmartone):

I guess that works then, great work! :)

OpenStudy (liv33):

Thanks for your help!

TheSmartOne (thesmartone):

It was my pleasure! :)